OpenStudy (gavin39):

the quantity of x minus 2, divided by 5 equals the quantity of 4x plus 2, divided by 25

OpenStudy (anonymous):

does the problem look like this \[\frac{ x-2 }{ 5 }=\frac{ 4x+2 }{ 25? }\]

OpenStudy (gavin39):

yes

OpenStudy (anonymous):

you are looking for the value of x?

OpenStudy (gavin39):

yeah

OpenStudy (anonymous):

if so start by multiplying both sides by 25 to rid yourself of the fraction you should end up with 5x-10=4x+2

OpenStudy (anonymous):

then you want to isolate the x variable on one side of the = bar so subtract the 4x from both sides you will end up with x-10=2

OpenStudy (anonymous):

and continue isolating the x you +10 to both sides leaving you with x=12

OpenStudy (anonymous):

did that help?